At 1000 K, Kp = 19.9 for the reaction Fe2O3(s)+3CO(g)⇌2Fe(s)+3CO2(g).

What are the equilibrium partial pressures of CO and CO2 if CO is the only gas present initially, at a partial pressure of 0.986 atm ?

Enter your answers numerically separated by a comma.
Answer #1

Sol.

Reaction :  

Fe2O3(s) + 3CO(g) <-----> 2Fe(s) + 3CO2(g)

initial 0.986 0

change - 3x + 3x

equilibrium 0.986 - 3x 3x

So , Kp = (PCO2)3 / (PCO)3  

19.9 = (3x)3 / (0.986 - 3x)3  

3x / (0.986 - 3x ) = (19.9)1/3 = 2.709  

3x = 2.709 ( 0.986 - 3x ) = 2.671 - 8.127x

x = 2.671 / 11.127 = 0.240

Therefore ,

Equilibrium partial Pressures are :

PCO = 0.986 - 3 × 0.240 =    0.266 atm    

PCO2 = 3 × 0.240 =    0.720 atm